A car starts from rest travelling with constant acceleration. if distance covered by it in 10th second of its journey is 19m, what will be the acceleration of the car-
 
1)4m/s^2
2)3m/s^2
3)2m/s^2
4)1m/s^2

Grade:11

1 Answers

Neeti
8 years ago
Formula for distance covered in nth second = S = u + a/2(2n-1)
In this question, u=0.
n = 10 , s= 19 . On applying values to the equation we get:
 
19 = a/2 (2x 10 -1 ) therefore a = 2m/s^2 
 
Please approve the answer if you think it’s useful :)